OpenStudy (pokemon23):

3(a-5)=2(2a+1) 3(a-5)=2(2a+1) @Mathematics

OpenStudy (pokemon23):

show all work?

OpenStudy (saifoo.khan):

3a - 15 = 4a + 2 3a - 4a = 15 + 2 -a = 17 a = -17

OpenStudy (pokemon23):

Saifoo.Khan does it again :)

jimthompson5910 (jim_thompson5910):

3(a-5)=2(2a+1) 3a-15=4a+2 3a=4a+2+15 3a-4a=2+15 -a=2+15 -a=17 a=17/(-1) a=-17

OpenStudy (saifoo.khan):

Like always! :D

OpenStudy (pokemon23):

nice Job

OpenStudy (saifoo.khan):

ty.

OpenStudy (anonymous):

3(a-5)= 2(2a+1). distribute 3 to "a" & 5. distribute 2 to 2a & 1. 3a-15= 4a+4 subtract four from both sides. 3a-19=4a. subtract 3a from both sides. -19=a. thats what I got. o: